OpenStudy (anonymous):

1154.65 m/s to km/hr

OpenStudy (shamim):

1 km=? m

OpenStudy (shamim):

1hr=?sec

OpenStudy (anonymous):

1000m

OpenStudy (anonymous):

3600s

OpenStudy (shamim):

1154.65m/s=(1154.65/1000)km/s Right?

OpenStudy (anonymous):

Yes

OpenStudy (shamim):

Because 1m=1/1000km

OpenStudy (shamim):

=(1154.65*3600/1000)km/hr

OpenStudy (shamim):

Because 1 s=1/3600 hr

OpenStudy (loser66):